OpenStudy (anonymous):

Find sin q if q is an angle in standard position and the point with coordinates (4, 3) lies on the terminal side of the angle. I don't understand what I'm suppose to be doing here.

OpenStudy (anonymous):

Basically you have this grid here:  So it just wants you to find angle X here, then evaluate it for sin x. (Assume q = x)

OpenStudy (anonymous):

So do I do sin(4)?

OpenStudy (anonymous):

No, you want the angle X first, remember that tan(X) = y/x = 3/4. So it follows that arctan or tan^-1(3/4) = 36.8698976 degrees. Then you just find sin(36.8698976) to get your answer.

OpenStudy (anonymous):

Oh okai I didn't know that, thank you c:

OpenStudy (anonymous):

No problem.